OpenStudy (anonymous):

Suppose that w and t vary inversely and that t = 1/5 one over five when w = 4. Write a function that models the inverse variation, and find t when w = 9.

OpenStudy (whpalmer4):

if w and t vary inversely, then w = k/t we know that t = 1/5 when w = 4, so we can plug those values into the equation and find the value of k. 4 = k/(1/5) 4 = 5k k = now we use the value we just found for k, along with the value w = 9, and find the value of t. we'll have to rearrange the equation to solve for t: w = k/t wt = k t = k/w plug in the values and solve